OpenStudy (bringmethehorizon_):

What is the value of x in the equation x − 7 = 11?

OpenStudy (bringmethehorizon_):

@GIL.ojei

OpenStudy (anonymous):

you have to take like terms to one side

OpenStudy (bringmethehorizon_):

4?

OpenStudy (anonymous):

\[x=7+11\]

OpenStudy (bringmethehorizon_):

18

OpenStudy (bringmethehorizon_):

7+11=18 ?

OpenStudy (anonymous):

no, when a figue or variable crosses the equality , its sign changes

OpenStudy (bringmethehorizon_):

Oohh. so subtraction ?

OpenStudy (anonymous):

so example, x-2=5 = x=5+2 else if it is x+2=5, to solve this cross 2 to meet 5 but change the sign x=5-2

OpenStudy (anonymous):

do u get that?

OpenStudy (bringmethehorizon_):

I'm still confused

OpenStudy (anonymous):

ok. lets solve this example, take the equality(=) to be a sign changer

OpenStudy (bringmethehorizon_):

if its 7-11 than its 4 ?

OpenStudy (anonymous):

look at x+4=2

OpenStudy (anonymous):

now, x is different from 4 but they are both on same side

OpenStudy (bringmethehorizon_):

The sign changes to -?

OpenStudy (anonymous):

but solve for x, you have carry 4 to the other side of 2 since they are same but when carrying it to the other side, the sign will change cos of the equality(=) separating the two sides

OpenStudy (bringmethehorizon_):

Since the 7 and the 11 are on the same side you have to carry it over and change the sign?

OpenStudy (anonymous):